Ask
your own question, for FREE!
Mathematics
14 Online
Help....;) Event A has a 0.3 probability of occuring and event B has a 0.4 probability of occuring. A and B are independent events. What is the probability that neither A or B occurs? 0.42 0.10 0.70 0.12
Still Need Help?
Join the QuestionCove community and study together with friends!
so....whats the answer from the choice?
so, 0.70?
haha yes of course : /
\[\large P(\bar {A})=1-0.3=0.7\] \[\large P(\bar{B})=1-0.4=0.6\] \[\large P(\bar{A} \cap \bar{B})=P(\bar{A}) \times P(\bar{B})=0.7\times0.6=you\ can\ calculate\]
you're the real mvp +kropot72 so, it's 0.42
Still Need Help?
Join the QuestionCove community and study together with friends!
Yes, 0.42 is correct.
The required probability is the intersection of the probabilities of A not occurring and B not occurring.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 day ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal